If the quantity is 0,1,2,3,4,5,6,7,8,9,10 then what will be the TFC,TC,AFC,AVC,ATC,MC if the tVC of the quantity is 0,25,40,50,60,80,110,150,300,500,900
1
Expert's answer
2020-03-17T09:24:49-0400

TFC is the same for every quantity. That's why TFC is not calculated.

MC is difference between TC and TC of quantity before. TFC in every TC the same, so MC is calculated as difference of VC divided by difference of quantity.

For 0 of quantity:

TC=TFC+0

Other indicators cannot be calculated for 0 of quantity

For 1 of quantity:

"TC=TFC+25"

"AFC=TFC\/1=TFC"

"AVC=25\/1=25"

"ATC=AFC+25"

"MC=(25-0)\/(1-0)=25"

For 2 of quantity:

"TC=TFC+40"

"AVC=40\/2=20"

"AFC=TFC\/2"

"ATC=AFC+20"

"MC=(40-25)\/(2-1)=15"

For 3 of quantity:

"TC=TFC+50"

"AVC=50\/3=16.66"

"AFC=TFC\/3"

"ATC=AFC+16.66"

"MC=(50-40)\/(3-2)=10"

For 4 of quantity:

"TC=TFC+60"

"AVC=60\/4=15"

"AFC=TFC\/4"

"ATC=AFC+15"

"MC=(60-50)\/(4-3)=10"

For 5 of quantity:

"TC=TFC+80"

"AVC=80\/5=16"

"AFC=TFC\/5"

"ATC=AFC+16"

"MC=(80-60)\/(5-4)=20"

For 6 of quantity:

"TC=TFC+110"

"AVC=110\/6=18.33"

"AFC=TFC\/6"

"ATC=AFC+18.33"

"MC=(110-80)\/(6-5)=30"

For 7 of quantity:

"TC=TFC+150"

"AVC=150\/7=21.42"

"AFC=TFC\/7"

"ATC=AFC+21.42"

"MC=(150-110)\/(7-6)"

For 8 of quantity:

"TC=TFC+300"

"AVC=300\/8=37.5"

"AFC=TFC\/8"

"ATC=AFC+37.5"

"MC=(300-150)\/(8-7)=150"

For 9 of quantity:

"TC=TFC+500"

"AVC=500\/9=55.55"

"AFC=TFC\/9"

"ATC=AFC+55.55"

"MC=(500-300)\/(9-8)=200"

For 10 of quantity:

"TC=TFC+900"

"AVC=900\/10=90"

"AFC=TFC\/10"

"ATC=AFC+90"

"MC=(900-500)\/(10-9)=400"